Dawit plans to consume two goods (good x and good y) with a limited budget. If Dawit’s budget line has intercepts 100 units of good x and 50 units of good y and price of good x (Px) is $100. Then, answer the next three questions. A) What are Dawit’s income and Price of good y (Py) B) What is the simplified version of Dawit’s budget line equation? C) If Dawit’s utility function is given by U(X; Y) = X0.5Y 0.5, how many units of good X and good Y would he consume if he choose the bundle that maximize his utility subject to his budget constraint?
1
Expert's answer
2021-09-19T18:35:00-0400

A) Dawit’s income is 100×100 = $10,000.

Price of good y is Py = 10,000/50 = $200.

B) The simplified version of Dawit’s budget line equation is:

100x + 200y = 10,000.

C) If Dawit’s utility function is given by U(X; Y) = X0.5Y0.5, then the bundle that maximize his utility subject to his budget constraint is at: "MUx\/MUy = Px\/Py."